# Anisotropy of quadratic forms over global fields of characteristic $\neq$ 2 is Diophantine

### v2

@@ 3.1. Notations and Some Examples
Let $\Delta_{1},\Delta_{2}$ be two finite subsets of $S_{K}^{f}$, then
(1) $\bigcap\limits_{\mathfrak{p}\in\Delta_{1}}\mathcal{O}_{\mathfrak{p}}+\bigcap\limits_{\mathfrak{p}\in\Delta_{2}}\mathcal{O}_{\mathfrak{p}}=\bigcap\limits_{\mathfrak{p}\in\Delta_{1}\cap\Delta_{2}}\mathcal{O}_{\mathfrak{p}}$,
(2) $\bigcap\limits_{\mathfrak{p}\in\Delta_{1}}\mathfrak{p}\mathcal{O}_{\mathfrak{p}}+\bigcap\limits_{\mathfrak{p}\in\Delta_{2}}\mathfrak{p}\mathcal{O}_{\mathfrak{p}}=\bigcap\limits_{\mathfrak{p}\in\Delta_{1}\cap\Delta_{2}}\mathfrak{p}\mathcal{O}_{\mathfrak{p}}$.
@@
For both (1) and (2), the left hand sides and obviously contained in the right hand sides. For the other direction, let $t=0$ or $1$. If $x\in K$ is an element such that $v_{\mathfrak{p}}(x)\geq t$ for $\mathfrak{p}\in\Delta_{1}\cap\Delta_{2}$, then by Chinese remainder theorem, there exists $x_{1}\in K$ such that $v_{\mathfrak{p}}(x_{1})\geq t,\forall\mathfrak{p}\in\Delta_{1}$ and $v_{\mathfrak{p}}(x_{1}-x)\geq t,\forall\mathfrak{p}\in\Delta_{2}\backslash\Delta_{1}$. Taking $x_{2}=x-x_{1}$, we have $x=x_{1}+x_{2}$ as the form of left hand sides. ∎
The following definition is a summary of Definition 3.21 and Definition 3.24 in [Par13], which are also written in [EM18] as Definition 3.13 and Definition 3.16.
Definition 3.8.
For each $\sigma\in\textup{Gal}(L/K)$, let
$\Phi_{\sigma}=\{p\in K^{\times}:(p)\in I_{\mathfrak{m}},\psi_{L/K}((p))=\sigma,\textup{ and }\mathbb{P}(p)\subset\mathbb{P}^{(1,1)}\cup\mathbb{P}^{\sigma}\};$
$\widetilde{\Phi_{\sigma}}:={K^{\times}}^{2}\cdot\Phi_{\sigma};$
$\Psi_{K}:=\Big\{(p,q)\in\widetilde{\Phi_{(1,1)}}\times\widetilde{\Phi_{(-1,-1)}}|\prod_{\mathfrak{p}|\mathfrak{m}}(ap,q)_{\mathfrak{p}}=-1\textup{ and }p\in a\cdot{K^{\times}}^{2}\cdot(1+J(R_{q}^{(-1,-1)}))\Big\}.$
For our convenience to use, we conclude some results in [EM18] and [Par13] as follows.
Proposition 3.9.
Let $K$ be a global field and keep notations above. The following holds:
  • (1)
    The sets $\Phi_{\sigma},\Psi_{K},R_{p}^{\sigma},J(R_{p}^{\sigma}),R_{p,q}^{(1,1)},J(R_{p,q}^{(1,1)})$ are all diophantine sets over $K$.
  • (2)
    For any $p\in\Phi_{\sigma}$ and $\sigma\in\textup{Gal}(L/K)$ with $\sigma\neq(1,1)$, $\mathbb{P}^{\sigma}(p)$ is nonempty.
  • (3)
    Let $\sigma\in\textup{Gal}(L/K)$ with $\sigma\neq(1,1)$, and let $\mathfrak{p}\nmid\mathfrak{m}$ be a prime of $K$ satisfying $\psi_{L/K}(\mathfrak{p})=\sigma$. Then there is an element $p\in\Phi_{\sigma}$ such that $\mathbb{P}^{\sigma}(p)=\{\mathfrak{p}\}$.
  • (4)
    For $(p,q)\in\Psi_{K}$, we have $\emptyset\neq\Delta_{ap,q}\cap\Delta_{bp,q}\subset I_{\mathfrak{m}}$.
  • (5)
    For each prime ideal $\mathfrak{p}_{0}$ satisfying $\mathfrak{p}_{0}\nmid\mathfrak{m}$ and $\psi_{L/K}(\mathfrak{p}_{0})=(1,1)$, there exists $(p,q)\in\Psi_{K}$ such that $\Delta_{ap,q}\cap\Delta_{bp,q}=\{\mathfrak{p}_{0}\}$.
@@
This is a conclusion of Lemma 3.14 and Lemma 3.17 in [EM18] for global function fields, and of Lemma 3.22 and Lemma 3.25 in [Par13] for number fields. ∎
Remark 3.10.
In fact, the sets $R_{p}^{\sigma},J(R_{p}^{\sigma}),R_{p,q}^{(1,1)},J(R_{p,q}^{(1,1)})$ are not only diophantine, but also in a uniform way rely on $p,q$. In other words, the sets
$\{(p,x)\in K:x\in R_{p}^{\sigma}\}$
$\{(p,x):x\in J(R_{p}^{\sigma})\}$
$\{(x,p,q):x\in R_{p,q}^{(1,1)}\}$
$\{(x,p,q):x\in J(R_{p,q}^{(1,1)})\}$
are diophantine. For these rings, the uniformity comes from the uniformity of $T_{a,b}$ (see Remark 3.2), Lemma 3.1 and Proposition 3.7(1). For these Jacobson ideals, the uniformity comes from Lemma 3.17 of [Par13] and Lemma 3.7(2).
Besides these properties above, the following lemma is implicit in [Par13] and [EM18] indeed, although not written as a lemma or theorem.
Lemma 3.11.
In (5) of Proposition 3.9, we can require that $v_{\mathfrak{p}_{0}}(p)=1,v_{\mathfrak{p}_{0}}(q)=0$ and $q$ is not a square modulo $\mathfrak{p}_{0}$.
Proof.
For the case of number fields, the proof of Lemma 3.25 (c) in [Par13] is made by a specific construction of $(p,q)$. When choosing $p,q\in K$, the condition (B) for $q$ is that the Legendre symbol $(\frac{q}{\mathfrak{p}_{0}})$ is equal to 1-1. This shows that $v_{\mathfrak{p}_{0}}(q)=0$ and $q$ is not a square modulo $\mathfrak{p}_{0}$. The condition (4) for $p$ is $(q,p)_{\mathfrak{p}_{0}}=-1$. By the calculation of Hilbert symbol, we have
$-1=(q,p)_{\mathfrak{p}_{0}}$ $=(-1)^{v_{\mathfrak{p}_{0}}(q)v_{\mathfrak{p}_{0}}(p){\frac{|\mathbb{F}_{\mathfrak{p}_{0}}|-1}{2}}}\big(\frac{q}{\mathfrak{p}_{0}}\big)^{v_{\mathfrak{p}_{0}}(p)}\big(\frac{p}{\mathfrak{p}_{0}}\big)^{v_{\mathfrak{p}_{0}}(q)}$
$=\big(\frac{q}{\mathfrak{p}_{0}}\big)^{v_{\mathfrak{p}_{0}}(p)},$
which implies that $v_{\mathfrak{p}_{0}}(p)$ is odd, which can be modified to $1$ after a free multiplication of a square in $K^{\times}$.
For the case of global function fields, it is in the proof of Lemma 3.17 (3) of [EM18]. In the construction, we also have the property that the Legendre symbol $(\frac{q}{\mathfrak{p}_{0}})$ is equal to 1-1. Since $\mathfrak{p}_{0}\in\Delta_{ap,q}$, we have $(ap,q)_{\mathfrak{p}_{0}}=-1$. By a calculation of the Hilbert symbol, and the fact that $\mathfrak{p}\nmid a$, we conclude $v_{\mathfrak{p}_{0}}(p)$ is odd, as stated in the third-to-last line of the proof of Theorem 1.3 in [EM18]. ∎
3.2. Diagonalization of quadratic forms
Let $f=\sum_{1\leq i,j\leq m}a_{ij}x_{i}x_{j}$ be a quadratic form of $m$ variables. We define the matrix $A=(A_{ij})$ by $A_{ii}=a_{ii}$ and $A_{ij}=\frac{a_{ij}+a_{ji}}{2}$ for $i\neq j$. Let $n$ be an integer from 0 to $m$.
$f$ is of rank $n$ if and only if there exists $(b_{i})_{1\leq i\leq n}$, in which $b_{i}$ are non-zero elements in $K$, and an invertible matrix $C:=(C_{rs})\in M_{m\times m}(K)$ such that $C^{t}AC=B$, where $B$ is a diagonal matrix of rank $n$: $\text{diag}(b_{1},...,b_{n},0,...,0)$.
For $2m^{2}$ variables $t_{ij},u_{rs}$, where $1\leq i,j,r,s\leq m$, we define $m^{2}$ polynomials $h_{kl},1\leq k,l\leq m$ as follows:
First we define the matrices of variables $T$ and $U$ by $T:=(t_{ij})_{1\leq i,j\leq m}$, $U:=(u_{rs})_{1\leq r,s\leq m}$. Then $U^{t}TU$ is a matrix whose entries are polynomials with variables $u_{rs}$ and $t_{ij}$. The polynomials $h_{kl}(t_{ij},u_{rs})$ are defined by the $(i,j)$-th entries of $U^{t}TU$. Moreover, the equation $C^{t}AC=B$ holds if and only if
  • $h_{k,l}(A_{ij},C_{ij})=0$ if $k\neq l$,
  • $b_{k}=h_{k,k}(A_{ij},C_{rs})$, if $1\leq k\leq n$
  • $h_{k,k}(A_{ij},C_{rs})=0$, if $n+1\leq k\leq m$.
In conclusion, we have the following property.
Proposition 3.12.
Let $f=\sum_{1\leq i,j\leq m}a_{ij}x_{i}x_{j}$ be a quadratic form in $m$ variables. Then there exist $m^{2}$ many polynomials $(h_{kl})_{1\leq k,l\leq m}\in\mathbb{Z}[t_{ij},u_{rs}]$ , depending only on the number of variables $m$, such that the following are equivalent:
(1) $f$ does not represent 0;
(2) There exist $b_{1},\dots,b_{m}\in K^{\times}$ such that the diagonal form $\sum_{i=1}^{m}b_{i}x_{i}^{2}$ does not represent 0, and the system of equations
$\begin{cases}h_{kl}(A_{ij},u_{rs})&=0,\quad 1\leq k\neq l\leq m\\ h_{kk}(A_{ij},u_{rs})&=b_{k},\quad 1\leq k\leq m\end{cases}$
admits a solution over $K$ in the variables $u_{rs}$.
Proof.
Let the polynomials $h_{kl}$ be the polynomials defined as above. Those polynomials only depends on $m$. If $f$ is of rank $m$, then the condition that there exist $b_{1},...,b_{m}\in K^{\times}$ such that the equation system admits a solution means exactly that $\sum_{i=1}^{m}b_{i}x_{i}^{2}$ is a diagonalization of $f=\sum_{1\leq i,j\leq m}a_{ij}x_{i}x_{j}$. Then $f$ does not represent 0 is equivalent to $\sum_{i=1}^{m}b_{i}x_{i}^{2}$ does not represent 0. If the rank of $f$ is less than $m$, then $f$ always represents 0. ∎
Under the condition that our main theorem (Theorem 1.2) holds, Corollary 1.4 can be obtained by using the diagonalization process (Proposition 3.12) above. Now we derive Corollary 1.4 from Theorem 1.2.
Proof.
Let $f=\sum_{1\leq i,j\leq m}a_{ij}x_{i}x_{j}$. By Proposition 3.12, the condition for $a_{ij}$ such that $f$ does not represents 0 is equivalent to the following condition:
$\exists b_{1},...,b_{m}\in K^{\times}$ and $C_{rs}\in K$ for $1\leq r,s\leq m$ such that
  • (1)
    $h_{k}(A_{ij};C_{rs})=b_{k}$ for all $1\leq k\leq m$,
  • (2)
    $\sum_{i=1}^{m}b_{i}x_{i}^{2}$ does not represent 0.
Then one can change the condition (2) to an equivalent diophantine condition by Theorem 1.2. ∎
@@ 4.1. Dealing with a Single Prime
Our strategy is to partition the set of primes into finitely many classes and handle the primes in each class in a uniform manner. As a first step, we need to consider the case of a single prime.
The following lemma, which is Lemma 4.4 in [EM18], is repeated here for completeness, as it constitutes a key ingredient in our proof.
@@
Let $\mathfrak{p}$ be a finite prime of $K$ with $|\mathbb{F}_{\mathfrak{p}}|$ odd. Fix $p,s\in K^{\times}$ so that $v_{\mathfrak{p}}(p)=1$ and $v_{\mathfrak{p}}(s)=0$ so that $\textup{red}_{\mathfrak{p}}(s)$ is not a square in the residue field of $\mathfrak{p}$. Then for $x,y\in K^{\times}$, $(x,y)_{\mathfrak{p}}=-1$ if and only if
@@
Besides, we will also need the following easy observation to obtain an equivalent form of Lemma 4.1.
@@
Let $\mathfrak{p}$ be a finite prime of $K$ with $|\mathbb{F_{\mathfrak{p}}}|$ odd. Fix $p,s\in K^{\times}$ so that $v_{\mathfrak{p}}(p)=1$ and $v_{\mathfrak{p}}(s)=0$ so that $\textup{red}_{\mathfrak{p}}(s)$ is not a square in the residue field of $\mathfrak{p}$. Then for $x,y\in K^{\times}$,
  • (1)
    $v_{\mathfrak{p}}(x)$ is even if and only if $x\in{K^{\times}}^{2}\cdot{\mathcal{O}_{\mathfrak{p}}^{\times}}$,
  • (2)
    $v_{\mathfrak{p}}(x)$ is odd if and only if $x\in p\cdot{K^{\times}}^{2}\cdot{\mathcal{O}_{\mathfrak{p}}^{\times}}$,
  • (3)
    $x$ is a square modulo $\mathfrak{p}$ if and only if $x\in{K^{\times}}^{2}\cdot(1+\mathfrak{p}\mathcal{O}_{\mathfrak{p}})$,
  • (4)
    $x$ is not a square modulo $\mathfrak{p}$ if and only if $x\in s\cdot{K^{\times}}^{2}\cdot(1+\mathfrak{p}\mathcal{O}_{\mathfrak{p}})$.
@@
(1): The valuation $v_{\mathfrak{p}}(x)$ is even if and only if the valuation can become 0 after a multiplication of a square, which is equivalent to $x\in{K^{\times}}^{2}\cdot\mathcal{O}_{\mathfrak{p}}^{\times}$.
(2) $v_{\mathfrak{p}}(x)$ is odd if and only if $v_{\mathfrak{p}}(x/p)$ is even.
(3) This is a restatement of $x$ is a square modulo $\mathfrak{p}$.
(4) This is due to the fact that the product of two quadratic nonresidue is a quadratic residue. ∎
@@
With the same notations as in Lemma 4.1, for $x,y\in K^{\times}$, $(x,y)_{\mathfrak{p}}=1$ if and only if
$((x\in{K^{\times}}^{2}\cdot\mathcal{O}_{\mathfrak{p}}^{\times})\vee(y\textup{ and }-xy\in{K^{\times}}^{2}\cdot(1+\mathfrak{p}\mathcal{O}_{\mathfrak{p}})))$
$\wedge((y\in{K^{\times}}^{2}\cdot\mathcal{O}_{\mathfrak{p}}^{\times})\vee(x\textup{ and }-xy\in{K^{\times}}^{2}\cdot(1+\mathfrak{p}\mathcal{O}_{\mathfrak{p}}))).$
@@
From Lemma 4.2 we can see that this lemma is a equivalent form of Lemma 4.1. ∎
@@
If $\mathfrak{p}\in S^{f}(K)$, then $f$ is isotropic over $K_{\mathfrak{p}}$ if and only if $a_{1}a_{2}a_{3}a_{4}\in K_{\mathfrak{p}}^{\times}$ and $(a_{1},a_{2})_{\mathfrak{p}}=-(-a_{3},-a_{4})_{\mathfrak{p}}$. So we just need to verify the following three kinds of sets are diophantine:
  • (1)

### v3

@@ 3.2. Diagonalization of quadratic forms
Let
$f=\sum_{1\leq i,j\leq m}a_{ij}x_{i}x_{j}$
be a quadratic form in $m$ variables over $K$. Define the symmetric matrix $A=[A_{ij}]_{1\leq i,j\leq m}$ by setting $A_{ii}=a_{ii}$ and $A_{ij}=\frac{a_{ij}+a_{ji}}{2}$ for $i\neq j$. Then there exist $m^{2}$ polynomials
$(h_{kl})_{1\leq k,l\leq m}\subset\mathbb{Z}[\overline{t_{ij}},\overline{u_{rs}}],$
depending only on $m$, where $\overline{t_{ij}}$ denotes the $m^{2}$ variables $(t_{ij})_{1\leq i,j\leq m}$ and $\overline{u_{rs}}$ denotes the $m^{2}$ variables $(u_{rs})_{1\leq r,s\leq m}$, such that the following statements are equivalent:
  • (1)
    The quadratic form $f$ is anisotropic over $K$.
  • (2)
    There exist $b_{1},\dots,b_{m}\in K^{\times}$ such that the diagonal quadratic form
    $\sum_{i=1}^{m}b_{i}x_{i}^{2}$
    is anisotropic over $K$, and there exist $(u_{rs})_{1\leq r,s\leq m}\in(K)^{m^{2}}$ satisfying
    $\begin{cases}h_{kl}(\overline{A_{ij}},\overline{u_{rs}})=0,&1\leq k\neq l\leq m,\\ h_{kk}(\overline{A_{ij}},\overline{u_{rs}})=b_{k},&1\leq k\leq m.\end{cases}$
@@
Consider the ring of $m\times m$ matrices with entries in $\mathbb{Z}[\overline{t_{ij}},\overline{u_{rs}}]$, denoted by
$M_{m\times m}(\mathbb{Z}[\overline{t_{ij}},\overline{u_{rs}}]).$
Define two matrices
$T\coloneqq[t_{ij}]_{1\leq i,j\leq m}\in M_{m\times m}(\mathbb{Z}[\overline{t_{ij}}])\subset M_{m\times m}(\mathbb{Z}[\overline{t_{ij}},\overline{u_{rs}}]),$
$U\coloneqq[u_{rs}]_{1\leq r,s\leq m}\in M_{m\times m}(\mathbb{Z}[\overline{u_{rs}}])\subset M_{m\times m}(\mathbb{Z}[\overline{t_{ij}},\overline{u_{rs}}]).$
Then
$U^{t}TU\in M_{m\times m}(\mathbb{Z}[\overline{t_{ij}},\overline{u_{rs}}]),$
and the $(k,l)$-th entry of $U^{t}TU$ is a polynomial
$h_{kl}\in\mathbb{Z}[\overline{t_{ij}},\overline{u_{rs}}].$
By construction, the polynomials $h_{kl}$ depend only on the number of variables $m$.
For matrices
$A\coloneqq[A_{ij}]_{1\leq i,j\leq m}\in M_{m\times m}(K),\qquad C\coloneqq[C_{rs}]_{1\leq r,s\leq m}\in M_{m\times m}(K),$
we define the evaluation of $H=(h_{kl})$ at $A$ and $C$ by
$H(A,C)_{kl}\coloneqq h_{kl}(\overline{A_{ij}},\overline{C_{rs}})\in K.$
This is well-defined since if $\textup{char}(K)=0$, then $\mathbb{Z}\subset K$, and if $\textup{char}(K)=p>0$, we interpret the polynomials modulo $p$.
Let $B=\text{diag}(b_{1},\dots,b_{m})$ be a diagonal matrix. Then the matrix equation
$C^{t}AC=B$
is equivalent to
$\begin{cases}h_{kl}(\overline{A_{ij}},\overline{C_{rs}})=0,&1\leq k\neq l\leq m,\\ h_{kk}(\overline{A_{ij}},\overline{C_{rs}})=b_{k},&1\leq k\leq m.\end{cases}$
If $f$ has full rank $m$, the existence of $b_{1},\dots,b_{m}\in K^{\times}$ and $C\in(K)^{m^{2}}$ satisfying this system is equivalent to the diagonalization of $f$ as
$\sum_{i=1}^{m}b_{i}x_{i}^{2}.$
Hence $f$ is isotropic over $K$ if and only if the diagonal form $\sum_{i=1}^{m}b_{i}x_{i}^{2}$ is isotropic over $K$.
If the rank of $f$ is less than $m$, then $f$ always isotropic, and the equivalence holds trivially.
In conclusion, condition (2) implies (1) by the above discussion, and condition (1) implies (2) since the symmetric matrix $A$ is diagonalizable over $K$. ∎
Under the assumption that our main theorem (Theorem 1.3) holds, Corollary 1.5 follows immediately from the diagonalization process in Proposition 3.11. We now give a precise derivation.
@@
Let $f=\sum_{1\leq i,j\leq m}a_{ij}x_{i}x_{j}$ be a quadratic form defined over $K$. Let the matrix $A=[A_{ij}]_{1\leq i,j\leq m}\in M_{m\times m}(K)$ be the matrix defined in Proposition 3.11. By Proposition 3.11, $f$ is isotrpoic over $K$ if and only if there exist $b_{1},\dots,b_{m}\in K^{\times}$ and $C_{rs}\in K$ ($1\leq r,s\leq m$) such that
$h_{kl}(\overline{A_{ij}},\overline{C_{rs}})$ $=0,$ $1\leq k\neq l\leq m,$
$h_{kk}(\overline{A_{ij}},\overline{C_{rs}})$ $=b_{k},$ $1\leq k\leq m,$
and the diagonal quadratic form $\sum_{i=1}^{m}b_{i}x_{i}^{2}$ is isotropic over $K$.
By Theorem 1.3, the condition that the diagonal form $\sum_{i=1}^{m}b_{i}x_{i}^{2}$ is isotropic over $K$ is diophantine. Therefore, the set of coefficients $(a_{ij})_{1\leq i,j\leq m}$ for which $f$ is anisotropic is diophantine, proving Corollary 1.5. ∎
@@ 4.1. Dealing with a Single Prime
Our strategy consists of partitioning the set of primes into finitely many classes and handling the primes in each class in a uniform manner. As a preliminary step, we consider the local case for a single prime.
The following lemma, which is a slightly different version of Lemma 4.4 in [EM18], is restated here for completeness, as it constitutes a key ingredient in our proof.
@@
Let $\mathfrak{p}$ be a finite prime of $K$ such that the residue field $\mathbb{F}_{\mathfrak{p}}$ has odd characteristic. Fix $p,s\in K^{\times}$ such that $v_{\mathfrak{p}}(p)$ is odd, $v_{\mathfrak{p}}(s)=0$ and the reduction $\textup{red}_{\mathfrak{p}}(s)$ is not a square in $\mathbb{F}_{\mathfrak{p}}$. Then for $x,y\in K^{\times}$, $(x,y)_{\mathfrak{p}}=-1$ if and only if
@@
Proof.
We modify the condition $v_{\mathfrak{p}}(p)=1$ used in Lemma 4.4 of [EM18] to the more general condition that $v_{\mathfrak{p}}(p)$ is odd. This modification does not alter the set $p\cdot{K^{\times}}^{2}\cdot\mathcal{O}_{\mathfrak{p}}^{\times}$. ∎
Besides, we also need the following elementary observation to obtain an equivalent form of Lemma 4.1, which is Lemma 4.3.
@@
Let $\mathfrak{p}$ be a finite prime of $K$ such that $|\mathbb{F}_{\mathfrak{p}}|$ is odd. Fix $p,s\in K^{\times}$ such that $v_{\mathfrak{p}}(p)=1$ and $v_{\mathfrak{p}}(s)=0$, with $\operatorname{red}_{\mathfrak{p}}(s)$ being a non-square in the residue field $\mathbb{F}_{\mathfrak{p}}$. Then for $x\in K^{\times}$, the following hold:
  1. (1)
    $v_{\mathfrak{p}}(x)$ is even if and only if $x\in{K^{\times}}^{2}\cdot\mathcal{O}_{\mathfrak{p}}^{\times}$.
  2. (2)
    $v_{\mathfrak{p}}(x)$ is odd if and only if $x\in p\cdot{K^{\times}}^{2}\cdot\mathcal{O}_{\mathfrak{p}}^{\times}$.
  3. (3)
    $x\in{K_{\mathfrak{p}}^{\times}}^{2}$ if and only if $x\in{K^{\times}}^{2}\cdot(1+\mathfrak{p}\mathcal{O}_{\mathfrak{p}})$.
  4. (4)
    $x\notin{K_{\mathfrak{p}}^{\times}}^{2}$ if and only if $x\in s\cdot{K^{\times}}^{2}\cdot(1+\mathfrak{p}\mathcal{O}_{\mathfrak{p}})\vee x\in p\cdot{K^{\times}}^{2}\cdot\mathcal{O}_{\mathfrak{p}}^{\times}$.
Consequently, we have the following two partitions of $K^{\times}$:
$K^{\times}$ $=({K^{\times}}^{2}\cdot\mathcal{O}_{\mathfrak{p}}^{\times})\sqcup(p\cdot{K^{\times}}^{2}\cdot\mathcal{O}_{\mathfrak{p}}^{\times})$
$=({K^{\times}}^{2}\cdot(1+\mathfrak{p}\mathcal{O}_{\mathfrak{p}}))\sqcup(s\cdot{K^{\times}}^{2}\cdot(1+\mathfrak{p}\mathcal{O}_{\mathfrak{p}}))\sqcup(p\cdot{K^{\times}}^{2}\cdot\mathcal{O}_{\mathfrak{p}}^{\times}).$
@@
(1) The valuation $v_{\mathfrak{p}}(x)$ is even if and only if it can be reduced to 0 by multiplication by a square in $K^{\times}$, which is equivalent to $x\in{K^{\times}}^{2}\cdot\mathcal{O}_{\mathfrak{p}}^{\times}$.
(2) First, if $u\in 1+\mathfrak{p}\mathcal{O}_{\mathfrak{p}}$, then $u\in{K_{\mathfrak{p}}^{\times}}^{2}$ by the local square theorem (see [O’M00], 63:1). Conversely, suppose $u\in{K_{\mathfrak{p}}^{\times}}^{2}$. By multiplying by a suitable square in $K^{\times}$, we may assume $v_{\mathfrak{p}}(u)=0$. Let $t\in K_{\mathfrak{p}}^{\times}$ be such that $u=t^{2}$. Since $K$ is dense in the local field $K_{\mathfrak{p}}$, there exists $\eta\in K$ such that $v_{\mathfrak{p}}(\eta-t)\geq 1$. Given that $v_{\mathfrak{p}}(t)=0$, we have $v_{\mathfrak{p}}(\eta)=\min\{v_{\mathfrak{p}}(\eta-t),v_{\mathfrak{p}}(t)\}=0$, which implies $\eta\in\mathcal{O}_{\mathfrak{p}}^{\times}$. It follows that
$u-\eta^{2}=(t-\eta)(t+\eta)\in\mathfrak{p}\mathcal{O}_{\mathfrak{p}},$
and therefore $u\in\eta^{2}+\mathfrak{p}\mathcal{O}_{\mathfrak{p}}=\eta^{2}(1+\mathfrak{p}\mathcal{O}_{\mathfrak{p}})$. This shows that $u\in{K^{\times}}^{2}\cdot(1+\mathfrak{p}\mathcal{O}_{\mathfrak{p}})$.
(4) For $x\notin{K_{\mathfrak{p}}^{\times}}^{2}$, we consider two cases based on the parity of the valuation:
  • If $v_{\mathfrak{p}}(x)$ is odd, then $x\in p\cdot{K^{\times}}^{2}\cdot\mathcal{O}_{\mathfrak{p}}^{\times}$ follows immediately from (2).
  • If $v_{\mathfrak{p}}(x)$ is even, we can write $x=p^{2k}u$ with $v_{\mathfrak{p}}(u)=0$ and $u\notin(K_{\mathfrak{p}}^{\times})^{2}$. According to [O’M00], 63:9, the index $\mathcal{O}_{\mathfrak{p}}^{\times}:(\mathcal{O}_{\mathfrak{p}}^{\times})^{2}]$ is $2$. Since $\operatorname{red}_{\mathfrak{p}}(s)$ is not a square in the residue field, $us$ must be a square in $K_{\mathfrak{p}}^{\times}$. By the result in (2), this implies $sx\in{K^{\times}}^{2}\cdot(1+\mathfrak{p}\mathcal{O}_{\mathfrak{p}})$, or equivalently, $x\in s\cdot{K^{\times}}^{2}\cdot(1+\mathfrak{p}\mathcal{O}_{\mathfrak{p}})$.
@@
With the same notation as in Lemma 4.1, for $x,y\in K^{\times}$, $(x,y)_{\mathfrak{p}}=1$ if and only if
$((x\in{K^{\times}}^{2}\cdot\mathcal{O}_{\mathfrak{p}}^{\times})\vee(y\textup{ and }-xy\in({K^{\times}}^{2}\cdot(1+\mathfrak{p}\mathcal{O}_{\mathfrak{p}}))\cup(p\cdot{K^{\times}}^{2}\cdot\mathcal{O}_{\mathfrak{p}}^{\times})))$
$\wedge((y\in{K^{\times}}^{2}\cdot\mathcal{O}_{\mathfrak{p}}^{\times})\vee(x\textup{ and }-xy\in({K^{\times}}^{2}\cdot(1+\mathfrak{p}\mathcal{O}_{\mathfrak{p}}))\cup(p\cdot{K^{\times}}^{2}\cdot\mathcal{O}_{\mathfrak{p}}^{\times}))).$
@@
By Lemma 4.1, $(x,y)_{\mathfrak{p}}=-1$ if and only if
To get $(x,y)_{\mathfrak{p}}=1$, we take the negation on this condition. By the partitions established in Lemma 4.2, the complement of the set $p\cdot{K^{\times}}^{2}\cdot\mathcal{O}_{\mathfrak{p}}^{\times}$ in $K^{\times}$ is ${K^{\times}}^{2}\cdot\mathcal{O}_{\mathfrak{p}}^{\times}$. Similarly, the complement of $s\cdot{K^{\times}}^{2}\cdot(1+\mathfrak{p}\mathcal{O}_{\mathfrak{p}})$ in the union of the two non-square classes is given by
$x\in({K^{\times}}^{2}\cdot(1+\mathfrak{p}\mathcal{O}_{\mathfrak{p}}))\cup(p\cdot{K^{\times}}^{2}\cdot\mathcal{O}_{\mathfrak{p}}^{\times}).$
Applying De Morgan’s laws and substituting these complements yields the desired characterization. ∎
In the dyadic case (which means $\mathfrak{p}|2$), one obtains a characterization of local squares similar to the one presented in Lemma 4.2 (3). However, due to the requirements of the stronger version of Hensel’s Lemma for even primes, a higher power of the prime ideal $\mathfrak{p}$ is necessary. Specifically, we have the following result:
Lemma 4.4.
Let $K$ be a number field and $\mathfrak{p}$ a prime dividing $2$. Let $e$ be the absolute ramification index. Then
${K^{\times}}^{2}\cdot(1+\mathfrak{p}^{2e+1}\mathcal{O}_{\mathfrak{p}})=K\cap(K_{\mathfrak{p}}^{\times})^{2}.$
Proof.
Suppose that $u\in 1+\mathfrak{p}^{2e+1}\mathcal{O}_{\mathfrak{p}}$. By 63:1 of [O’M00], $u$ is a local square, so $u\in{K_{\mathfrak{p}}^{\times}}^{2}$. Since $u\in K$ by assumption, it follows that $u\in K\cap{K_{\mathfrak{p}}^{\times}}^{2}$. This establishes the inclusion ${K^{\times}}^{2}\cdot(1+\mathfrak{p}^{2e+1}\mathcal{O}_{\mathfrak{p}})\subseteq K\cap{K_{\mathfrak{p}}^{\times}}^{2}$.
Conversely, let $x\in K\cap{K_{\mathfrak{p}}^{\times}}^{2}$. We can write $x=\pi^{2k}u$, where $v_{\mathfrak{p}}(\pi)=1$ and $u\in\mathcal{O}_{\mathfrak{p}}^{\times}\cap{K_{\mathfrak{p}}^{\times}}^{2}$. Let $s\in\mathcal{O}_{\mathfrak{p}}^{\times}$ be such that $u=s^{2}$. If $s\in K$, the result is trivial. If $s\notin K$, then by the density of $K$ in $K_{\mathfrak{p}}$, there exists $\eta\in K$ such that
$v_{\mathfrak{p}}(\eta-s)\geq 2e+1.$
Since $v_{\mathfrak{p}}(s)=0$ and $2e+1>0$, the properties of valuations imply $v_{\mathfrak{p}}(\eta)=v_{\mathfrak{p}}(\eta-s+s)=\min\{v_{\mathfrak{p}}(\eta-s),v_{\mathfrak{p}}(s)\}=0$, hence $\eta\in\mathcal{O}_{\mathfrak{p}}^{\times}$.
To show that $u\eta^{-2}\in 1+\mathfrak{p}^{2e+1}\mathcal{O}_{\mathfrak{p}}$, we consider the difference:
$u-\eta^{2}=(s-\eta)(s+\eta).$
By our choice of $\eta$, we have $v_{\mathfrak{p}}(s-\eta)\geq 2e+1$. For the second factor, since $s+\eta=(s-\eta)+2\eta$, and noting that $v_{\mathfrak{p}}(s-\eta)\geq 2e+1>e=v_{\mathfrak{p}}(2\eta)$, we have
$v_{\mathfrak{p}}(s+\eta)=v_{\mathfrak{p}}(2\eta)=e.$
Therefore, $v_{\mathfrak{p}}(u-\eta^{2})=v_{\mathfrak{p}}(s-\eta)+v_{\mathfrak{p}}(s+\eta)\geq(2e+1)+e=3e+1$. In particular, this implies $v_{\mathfrak{p}}(u-\eta^{2})\geq 2e+1$. Since $v_{\mathfrak{p}}(\eta^{2})=0$, we conclude that
$u\in\eta^{2}(1+\mathfrak{p}^{2e+1}\mathcal{O}_{\mathfrak{p}}),$
so $x\in{K^{\times}}^{2}\cdot(1+\mathfrak{p}^{2e+1}\mathcal{O}_{\mathfrak{p}})$. This proves the reverse inclusion $K\cap{K_{\mathfrak{p}}^{\times}}^{2}\subseteq{K^{\times}}^{2}\cdot(1+\mathfrak{p}^{2e+1}\mathcal{O}_{\mathfrak{p}})$. ∎
@@
If $\mathfrak{p}\in S^{f}(K)$, then by Theorem 2.3(3), $f$ is anisotropic over $K_{\mathfrak{p}}$ if and only if $a_{1}a_{2}a_{3}a_{4}\in{K_{\mathfrak{p}}^{\times}}^{2}$ and $(a_{1},a_{2})_{\mathfrak{p}}=-(-a_{3},-a_{4})_{\mathfrak{p}}$. To establish that this condition defines a diophantine set, it suffices to verify that the following three types of sets are diophantine:
  1. (1)